1. Present Value and Multiple Cash Flows [LO1] Investment X offers to pay you $4,200 per year for eight years, whereas Investment Y offers to pay you $6,100 per year for five years. Which of these cash flow streams has the higher present value if the discount rate is 5 percent? If the discount rate is 15 percent?

2. Future Value and Multiple Cash Flows [LO1] Fuente, Inc., has identified an investment project with the following cash flows. If the discount rate is 8 percent, what is the future value of these cash flows in Year 4? What is the future value at a discount rate of 11 percent? At 24 percent?

Year

Cash Flow

1

$1,075

2

1,210

3

1,340

4

1,420
Answer #1

1.
a)
Present value if the discount rate is 5 percent:

Investment X:

PMT = 4200
Nper = 8
Rate = 5%
FV = 0

Present value can be calculated by using the following excel formula:
=PV(rate,nper,pmt,fv)
=PV(5%,8,-4200,0)
= $27,145.49

Investment Y:

PMT = 6100
Nper = 5
Rate = 5%
FV = 0

Present value can be calculated by using the following excel formula:
=PV(rate,nper,pmt,fv)
=PV(5%,5,-6100,0)
= $26,409.81

Investment X has the higher present value if the discount rate is 5%.


Present value if the discount rate is 15 percent:

Investment X:

PMT = 4200
Nper = 8
Rate = 15%
FV = 0

Present value can be calculated by using the following excel formula:
=PV(rate,nper,pmt,fv)
=PV(15%,8,-4200,0)
= $18,846.75

Investment Y:

PMT = 6100
Nper = 5
Rate = 15%
FV = 0

Present value can be calculated by using the following excel formula:
=PV(rate,nper,pmt,fv)
=PV(15%,5,-6100,0)
= $20,448.15

Investment Y has the higher present value if the discount rate is 15%

Note: As per HOMEWORKLIB RULES, only first question can be answered. Post the rest of the questions separately.